# Privacy in Speech Technology

### v2

@@ IX Discussion and Conclusions
This paper is a tutorial on privacy for speech technology. Its most notable novelty is an exhaustive categorization of threats (see figs. 1 and III). Protections against those threats are further categorized according to whether they relate to the private message or side information. The pertinent difference is that transmitting a private message is the whole purpose of communication, and there is not much we can do to protect it other than encryption. With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections. The primary approach is, however, to remove as much of the side information as possible as early as possible. As the private message is all that communication we need, all side-information should be removed to the extent possible. Such removal rapidly, however, demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message. It is thus not always clear what constitutes a legitimate private message.
@@
Second, we find that privacy and ethics are largely overlapping challenges. Our ethical values govern our preferences for privacy. Most potential breaches of ethics in speech technology are based on breaching privacy. That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy. Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [heikkila2023cambridge_ai].
@@
While management of acquiring informed consent has established traditions and best practices for most interface types [friedman2000informed], speech, audio, and ambient systems are notably unique. Namely, acoustic information is a time-varying stream. Reading out a pages-long consent form before an interaction can start is clearly much too obtrusive and unnecessarily detailed. Privacy requirements also vary over time. Consent should thus be acquired on an actual need basis. In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is actually needed.
@@
The metrics discussed in section V are all related to specific categories of private information and in particular, we can provide protection only to identified threats. For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background. We thus need methods for evaluating privacy jointly with respect to all categories of private information except the private message.
@@
Empirical metrics are generally based on a model of the signal or the attacker. The metrics are thus subject to change when those models are improved in the future, and it will likely expose new threats. Though it is likely difficult, it would be extremely useful if we could characterize, for example, privacy threats as a function of computational complexity, similarly as cryptography provides assurances over security based on the difficulty of solving certain difficult computational problems. Where cryptographic security is often seen as a threshold of computational complexity, however, for privacy in speech applications, it would be beneficial to also assess the leak magnitude (e.g. in bits) as a function of computational complexity. Over time, improvements in efficiency reduce algorithmic complexity, and as hardware with better capabilities becomes available, this would allow characterizing how protections will survive the test of time.
@@
The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [roberts2007biometric]. Characterization of speech-specific attacks has received much less attention, though figs. 1 and III in this paper are attempts in that direction. Further work is needed to gain privacy-protections on a systems level.
@@
Privacy research is categorically focused on personal and user-centric privacy. However, speech is, by definition, communication between multiple agents, and when exposed, threatens all participants simultaneously. This is not an issue from a legal point of view, because privacy protections apply to all individual users equally. However, from an authorization and consent management perspective, this is an underappreciated issue. If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording. Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree. We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
If we could disentangle all categories of speech information as in fig. 14, it would be easy to anonymize each category to an appropriate degree. This approach thus seemingly solves all our problems. The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement. The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined and there is significant overlap between them. We cannot even demonstrate that this would be a complete list of information categories. Without exact definitions of those categories, we have no hope of developing methods for them. An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to explicitly define categories. The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories. Still, since disentanglement is the ideal solution, that should continue to be a central focus of research.
@@
Most of the speech-specific research on privacy has focused on privacy-preserving processing and systems structures. This is useful because it is the mandatory prerequisite for privacy-preserving technology. However, as discussed in section VI-B, users’ experience of services is to some extent independent of the objective level of privacy. We need many more user studies on how, for example, voice characteristics and word choices influence trust, how the privacy level can be monitored during interactions and how changes are notified, how the environment and content of interaction influence user experiences, etc. By improving the user experience with respect to privacy, we are likely to improve user satisfaction and retention of the overall service, while also improving the service objectively.
In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels. While further improvements in smart technology are expected to improve the utility of the technology, it likely also introduce new threats. The protection of privacy in speech technology has thus been important already for a long time and the importance is increasing. Fortunately, research in the area has picked up speed and this tutorial presents the most important concepts, approaches, and methodology. It is, however, likely that fundamental results and new technologies will be introduced in the near future. This is thus an exciting time for researchers in the area.
@@ Acknowledgments
A majority of the content in this paper has risen directly from discussions with a long list of colleagues, who were far too many to include as co-authors. These colleagues include, but are not limited to, in alphabetical order, Chris Brzuska, Mads Græsbøll Christensen, Sneha Das, Nicholas Evans, Johannes Fischer, Francisco Florez, Florin Ghido, Meeri Haataja, Ivan Habernal, Aki Härmä, Dorothea Kolossa, Michael Laakasuo, Martha Larson, Anna Leschanowsky, Rainer Martin, Joachim Meyer, Sebastian Möller, Andreas Nautsch, Birgit Popp, Nitin Sawhney, Ingo Siegert, Stephan Sigg, Rech Silas, Isabel Trancoso, Sanna Toropainen, Emmanuel Vincent, Jennifer Williams, and Pablo Pérez Zarazaga. I will remain forever grateful.
[…]
[Uncaptioned image] Tom Bäckström D.Sc. (tech.) is an associate professor at Aalto University, Finland (2019-). He obtained his Master’s and Doctoral degrees at Helsinki University of Technology (the predecessor of Aalto) in 2001 and 2004, respectively. During his time at International Audio Laboratories Erlangen, Germany (2008-2016), he made contributions to several international speech and audio coding standards such as MPEG USAC and 3GPP EVS, and became a professor (W2) at Friedrich-Alexander University Erlangen-Nürnberg (FAU) (2012-2016). Before his current position, he was a professor practice at Aalto University (2016-2019). He was the initiator and chair of ISCA SIG “Security and Privacy in Speech Communication” (2019-2022). His current research interests include privacy, coding, enhancement, and transmission of speech as well as machine learning.

### v3

@@ IX Discussion and conclusions
This paper is a tutorial on privacy for speech technology. Its most notable novelty is an exhaustive categorization of threats (see figs. 1 and IV). Protections against those threats are further categorized according to whether they relate to the target information or side information. The pertinent difference is that transmitting target information is the whole purpose of communication, and there is not much we can do to protect it other than encryption. With side-information, that is, all the other information that is bundled into a speech, like health status and gender identity, we have a much larger arsenal of protections. The primary approach is, however, to remove as much of the side information as possible as early as possible. As the target information is all the communication we need, all side-information should be removed to the extent possible. Such removal, however, rapidly demonstrates that paralinguistic information like speech style is often very useful in conveying the intended message. It is thus not always clear what constitutes the legitimate, target information.
@@
Second, we find that privacy and ethics are largely overlapping challenges. Our ethical values govern our preferences for privacy. Most potential breaches of ethics in speech technology are based on breaching privacy. That means that we need a society-wide ethical discussion about what is allowed with respect to user privacy. Such discussions are needed to prevent a Cambridge Analytica-style scandal for speech technology [219].
@@
While management of acquiring informed consent has established traditions and best practices for most interface types [236], speech, audio, and ambient systems are notably unique. Namely, acoustic information is a time-varying stream. Reading out a pages-long consent form before an interaction can start is clearly much too obtrusive and unnecessarily detailed. Privacy requirements also vary over time. Consent should thus be acquired on an actual need basis. In addition to being more usable, it would also make choices better connected to the actual needs since consent is acquired only once it is needed.
@@
The metrics discussed in section VI are all related to specific categories of private information and in particular, we can provide protection only to identified threats. For example, we can measure the threat to privacy related to health information, but that does not say anything about the threat related to information about ethnic background. We thus need methods for evaluating privacy jointly with respect to all categories of private information except the target information.
@@
Empirical metrics are generally based on a model of the signal or the attacker. The metrics are thus subject to change when those models are improved in the future, and it will likely expose new threats. Though it is likely difficult, it would be extremely useful if we could characterize, for example, privacy threats as a function of computational complexity, similarly to how cryptography provides assurances over security based on the difficulty of solving certain difficult computational problems. Where cryptographic security is often seen as a threshold of computational complexity, however, for privacy in speech applications, it would be beneficial to also assess the leak magnitude (e.g., in bits) as a function of computational complexity. Over time, efficiency improvements reduce algorithmic complexity, and as hardware with better capabilities becomes available, this would allow us to characterize how protections will survive the test of time.
@@
The biometrics community has a long tradition of characterizing attacks as, for example, opportunistic and motivated attacks, as well as in terms of the potential rewards of such attacks [237]. Characterization of speech-specific attacks has received much less attention, though figs. 1 and IV in this paper are attempts in that direction. Further work is needed to gain privacy-protections on a systems level.
@@
Privacy research is categorically focused on personal and user-centric privacy. However, speech is, by definition, communication between multiple agents, and when exposed, threatens all participants simultaneously. This is not an issue from a legal point of view, because privacy protections apply to all individual users equally. However, from an authorization and consent management perspective, this is an underappreciated issue. If user A records a discussion with user B, then both clearly have some level of ownership and privacy requirements on that recording. Another case is smart technology with multiple users, like smart TVs; even if one user has consented to data collection, that does not mean that others would agree [222]. We do not yet have any widely accepted standard approaches for handling privacy, ownership, and consent in such multi-user scenarios.
@@
If we could disentangle all categories of speech information as in fig. 14, it would be easy to anonymize each category to an appropriate degree. This approach thus seemingly solves all our problems. The issue is that we do not yet have sufficiently sophisticated methods for such disentanglement. The difficulty in developing disentanglement algorithms is that information categories in table I are vaguely and heuristically defined and there is significant overlap between them. We cannot even demonstrate that this would be a complete list of information categories. Without exact definitions of those categories, we have no hope of developing methods for them. An alternative approach is to use representation learning methods to create unsupervised clustering of information categories, thus circumventing the need to define categories explicitly. The compromise is that we cannot guarantee that the learned representations correspond to heuristically meaningful categories. Still, since disentanglement is the ideal solution, that should continue to be a central focus of research.
@@
Most of the speech-specific research on privacy has focused on privacy-preserving processing and system structures. This is useful because it is the mandatory prerequisite for privacy-preserving technology. However, as discussed in section VII-C, users’ experience of services is to some extent independent of the objective level of privacy. We need many more user studies on how, for example, voice characteristics and word choices influence trust, how the privacy level can be monitored during interactions and how changes are notified, how the environment and content of interaction influence user experiences, etc. By improving the user experience with respect to privacy, we are likely to improve user satisfaction and retention of the overall service, while also improving the service objectively.
In conclusion, threats and breaches of privacy have significant negative consequences on individual, societal, ethical, and economic levels. While further improvements in smart technology are expected to improve the utility of the technology, it will likely also introduce new threats. The protection of privacy in speech technology has thus been important already for a long time, and its importance is increasing. Fortunately, research in the area has picked up speed, and this tutorial presents the most important concepts, approaches, and methodology. It is, however, likely that fundamental results and new technologies will be introduced in the near future. This is thus an exciting time for researchers in the area.
@@ Acknowledgments
A majority of the content in this paper has risen directly from discussions with a long list of colleagues, who were far too many to include as co-authors. These colleagues include, but are not limited to, in alphabetical order, Chris Brzuska, Mads Græsbøll Christensen, Sneha Das, Nicholas Evans, Johannes Fischer, Francisco Florez, Florin Ghido, Meeri Haataja, Ivan Habernal, Aki Härmä, Dorothea Kolossa, Michael Laakasuo, Martha Larson, Anna Leschanowsky, Rainer Martin, Joachim Meyer, Sebastian Möller, Andreas Nautsch, Francesco Nespoli, Birgit Popp, Silas Rech, Nitin Sawhney, Ingo Siegert, Stephan Sigg, Francisco Teixeira, Isabel Trancoso, Sanna Toropainen, Ville Vestman, Emmanuel Vincent, Jennifer Williams, and Pablo Pérez Zarazaga. I will remain forever grateful.
A part of this work was supported by “Designing Inclusive & Trustworthy Digital Public Services for Migrants in Finland (Trust-M)” (Grant Number: 353529) funded by the Strategic Research Council of Finland.